OpenStudy (anonymous):

The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers?

OpenStudy (phi):

This is algebra. Give the two unknown numbers names. For example, x and y now use x and y in two different equations: the sum of x and y is 110. Can you make that into an equation? the difference of x and y is 28. Can you make that into an equation?

OpenStudy (experimentx):

x+y = 110 x-y = 28 x = 69 y = 41 seems only one possible solution

OpenStudy (anonymous):

x+y= 110 and x-y=28

OpenStudy (anonymous):

Thank you so much(:!

OpenStudy (phi):

Yes. Good job. One easy way to solve is add the two equations together x+y + (x-y) = 110+28 the y's cancel, and the x's add up to 2x, so you get 2x= 138 x= 69 because x+y= 110 we know 69+y= 110 subtract 69 from both sides to get y= 41

OpenStudy (anonymous):

Thank you!! This makes so much more sense to me now!